OpenStudy (juan1857):

Find the polar coordinates of each point with the given rectangular coordinates. Use 00. (0,1)

jimthompson5910 (jim_thompson5910):

Given point is (x,y) = (0,1) Use the formulas \[\Large r = \sqrt{x^2+y^2}\] \[\Large \theta = \arctan\left(\frac{y}{x}\right)\] basically the same as before but instead of (a,b) I'm using (x,y)

jimthompson5910 (jim_thompson5910):

because x = 0 this means that y/x will be undefined. So the result of arctangent will be either pi/2 or 3pi/2 where tangent is undefined In this case, (0,1) is above the x axis so the angle is theta = pi/2

OpenStudy (juan1857):

oh ok so when x is undefined, theta =pi/2 ?

jimthompson5910 (jim_thompson5910):

if x = 0 and y > 0, then theta = pi/2 if x = 0 and y < 0, then theta = 3pi/2

OpenStudy (juan1857):

ahh Ok, And @jim_thompson5910 does it again!!!!! Thanks again.

jimthompson5910 (jim_thompson5910):

no problem